OpenStudy (anonymous):

Using double angle identities: Find sin (2x). if tan x =12/5 and pi < x < 3pi/2

OpenStudy (anonymous):

tan x = opp/adj. Since they're in pi <x< 3pi/2, they are both negative. So: opp= -12 adj = -5 Use (-12)^2 + (-5)^2 = c^2 . hypotenuse = c= 13 , hyp has no negative sign. sin u = 2sin u cos u sin (2x) = 2 (-12/13) (-5/13) = 120/169 = 0.7101

OpenStudy (anonymous):

@study100 thank you